OpenStudy (anonymous):

Hey! I'm stuck on some problems on my homework. 1. Find the following limit. lim as x approaches infinity of (square root of 36x squared +x -6x) note: -6x is outside the square root.

OpenStudy (anonymous):

\[\lim_{x\rightarrow\infty}{[\sqrt{3x^2+x}-6x]}\] Is that it?

OpenStudy (anonymous):

yes. i don't know how to write it like that lol

OpenStudy (anonymous):

Haha its the code of LaTex. Learn it then practice using it because it can simplify things a lot and write equations with any complexity. For example: \[y=\left[(1-n)\frac{\int{e^{(1-n)\int{p(x)dx}}q(x)}+c_1}{e^{(1-n)\int{p(x)dx}}}\right]^{\frac{1}{1-n}}\] Lol so its pretty neat. Anyways. back to the question. I haven't even began simplifying anything, but graphically, it appears: \[\lim_{x\rightarrow\infty}{\sqrt{3x^2+x}-6x}=-\infty\]

OpenStudy (anonymous):

Where is the LaTex code located? and how did you get -infinity ??

OpenStudy (anonymous):

Well I mean, some sites worth checking out are: http://www.latex-project.org/guides/ http://www.mathjax.org/resources/docsindex/ And I said just by graphing it, it looks like that

OpenStudy (anonymous):

ok thanks alot :)